7.) A businessman is faced with the prospect of fluctuating future budget for the maintenance of the generator. During the first 5 years, P1,000 per year will be budgeted. During the second five years, the annual budget will be P1,500 per year. In addition, P3,500 will be budgeted for an overhaul of the machine at the end of the fourth year and another P3,500 for an overhaul at the end of 8th year. Assuming compound interest at 6% per annum, what is the equivalent annual cost of maintenance?


Ans. P1,888.87


1
Expert's answer
2022-03-25T15:06:58-0400

"Annuity factor=\frac{1-{1 \over (1+r)^t}}{r}=\frac{1-{1 \over (1+0.06)^10}}{0.06}=15.215"

"EAC={NPV \over Annuity factor}"

"NPV={1000 \over 1+0.06}+{1000 \over (1+0.06)^2}+{1000 \over (1+0.06)^3}+{1000+3500 \over (1+0.06)^4}+{1000 \over (1+0.06)^5}+{1500 \over (1+0.06)^6}+{1500 \over (1+0.06)^7}+{1500+3500 \over (1+0.06)^8}+{1500 \over (1+0.06)^9}+{1500 \over (1+0.06)^10}=13064.627"

"EAC={13064.627 \over 15.215}=858.664"